OpenStudy (oatmeal):

(1/2)2 how do you do this?

OpenStudy (lalaly):

\[\frac{1}{2} \times 2 = \frac{2}{2} =1\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

1/2 + 1/2 = 1

OpenStudy (amistre64):

\[(\frac{1}{2})^2=\frac{1}{2}*\frac{1}{2}=\frac{1}{4}\]